10. Julia plays tennis. She can score a point from the base line with a probability 0.55. How many times does she need to hit the ball so that the probability of Julia Scoring a point exceeds 0.999 ?
Answer #1

P(scoring a point) = 0.55

P(not scoring a point) = 1 - 0.55 = 0.45

Let n be the number of times Julia need to hit the ball so that the probability of she scoring a point exceeds 0.999

0.999 = P(she scoring at least one point)

0.999 = 1 - P(she not scoring on any hits)

0.999 = 1 - 0.45n

0.45n = 0.001

n log 0.45 = log 0.001

n = log 0.001 / log 0.45

n = 8.65

\approx 9